# Birds look like cars: Adversarial analysis of intrinsically interpretable deep learning

### v1

@@ 1 $\mathbf{Z},\mathbf{P},\mathbf{I},\hat{\theta}_{g}\leftarrow[\;],[\;],[\;],[\;]$
In what follows, let $\mathcal{D}_{\mathrm{OOD}}$ denote a set of inputs an adversary chooses for prototype manipulation. The naive approach is to only substitute $\mathcal{D}_{\mathrm{train}}$ with $\mathcal{D}_{\mathrm{OOD}}$ for the single prototype projection stage during training. Specifically for ProtoViT, we propose to randomly assign inputs in $\mathcal{D}_{\mathrm{OOD}}$ to the original labels from $\mathcal{Y}$ to satisfy the additional constraint mentioned above. We found this baseline to achieve subpar results (as later shown in Table 1) and thus implemented a simple matching algorithm instead. We substitute latent representations stored in $\theta_{g}$ with the closest embeddings of parts of inputs in $\mathcal{D}_{\mathrm{OOD}}$ as denoted with $\hat{\theta}_{g}$ and given in Algorithm 1. In practice, iterating over the prototype projection set (Alg. 1, line 2) is executed in batches similar to training. Note that prototype substitution can be done at any point, e.g. after the first stage of training encoder parameters $\theta_{f}$, or after the last stage of training classification parameters $\theta_{h}$ when parameters $\theta_{f}$ are frozen. In general, substituting prototypes introduces an approximation error $||\theta_{g}-\hat{\theta}_{g}||$ that propagates to errors in predictions of $g$ and finally $h$. One can correct the predictions of $h$ by fine-tuning parameters $\theta_{h}$, which we further analyze in Section 4.1.
@@
Backdoor attacks.
The goal of a backdoor attack is to change the model’s prediction for inputs with a specific trigger. Such an attack could be discovered with post-hoc explanation methods (Noppel et al., 2023). Thus, a successful disguising backdoor attack in the context of interpretability assumes concealing the trigger in a visualization of the model’s explanation. Figure 2 shows an intuition behind our attack on interpretable deep learning models. In an extreme scenario, an adversary might aim to exaggerate the evidence for a manipulated prediction so as to distract from the attack, aka red herring.
Note that natural semantic triggers already exist in data in the form of bias, e.g. commonly in medical images (Nauta et al., 2023a). Figure 3 shows features occurring coincidentally with the predicted outcome while having no causal influence on it (see Mikołajczyk et al., 2022, and references given there for a discussion on this issue). Although we implement the trigger as a small adversarial patch in an image corner to observe the qualitative result, our method is general in nature. Intuitively, such a backdoor could work because there exists no prototype to be matched with a trigger, so the model chooses another prototype to match instead. However, as we show in Section 4.1, the model is able to learn superficial prototypes with Algorithm 1 if needed.
@@
To embed our backdoor, we fine-tune a trained model $\theta$ on an extended dataset $\{\mathcal{D}_{\mathrm{train}},\mathcal{D}_{\mathrm{trigger}}\}$ where $\mathcal{D}_{\mathrm{trigger}}\coloneqq\left\{\left(\mathbf{x}^{(n+1)},y^{(n+1)}\right),\ldots,\left(\mathbf{x}^{(n+m)},y^{(n+m)}\right)\right\}\in\mathcal{X}_{\mathrm{trigger}}\times\mathcal{Y}$ with $m$ new inputs that include the trigger and a changed class label. For example, one can switch the class $y^{(n+i)}\coloneqq 1-y^{(i)}$ for an input $\mathbf{x}^{(n+i)}\coloneqq\mathbf{x}^{(i)}\mathbin{\mathchoice{\ooalign{$\displaystyle\circ$\cr$\displaystyle\cdot$\cr}}{\ooalign{$\textstyle\circ$\cr$\textstyle\cdot$\cr}}{\ooalign{$\scriptstyle\circ$\cr$\scriptstyle\cdot$\cr}}{\ooalign{$\scriptscriptstyle\circ$\cr$\scriptscriptstyle\cdot$\cr}}}\mathbf{t}$ containing an adversarial patch $\mathbf{t}$. In general, we aim to align embeddings (interpretations) between $\mathbf{x}^{(n+i)}$ and $\mathbf{x}^{(i)}$ while the model learns to predict for them different classes. While this could be implemented for any model architecture, we conveniently observe that the original loss function in PIP-Net can already be modified to achieve our goal.
We define the loss function for adversarial fine-tuning as follows:
$\mathcal{L}_{\mathrm{adv}}(\hat{\theta})=\lambda_{\mathrm{C}}\mathcal{L}_{\mathrm{C}}+\lambda_{\mathrm{A},\neg\mathrm{T}}\mathcal{L}_{\mathrm{A},\neg\mathrm{T}}+\lambda_{\mathrm{A},\mathrm{T}}\mathcal{L}_{\mathrm{A},\mathrm{T}}+\lambda_{\mathrm{U},\neg\mathrm{T}}\mathcal{L}_{\mathrm{U},\neg\mathrm{T}}+\lambda_{\mathrm{U},\mathrm{T}}\mathcal{L}_{\mathrm{U},\mathrm{T}}$ (2)
where $\mathrm{T}$ denotes operations on embeddings of inputs from $\mathcal{D}_{\mathrm{trigger}}$ and $\neg\mathrm{T}$ on the original ones. Algorithm 2 describes in detail the exemplary computation of a backdoor attack for binary classification with trigger $\mathbf{t}$, where iterating over the training set (Alg. 2, line 3) is executed in batches in practice. Our implementation greatly benefits from using 2 GPUs to store model parameters, inputs, and outputs between $\theta$ and $\hat{\theta}$. Manipulating $\lambda_{\mathrm{A},\neg\mathrm{T}}$ and $\lambda_{\mathrm{A},\mathrm{T}}$ hyperparameters produces a spectrum of results between disguising and red herring backdoor attacks. For example, setting $\lambda_{\mathrm{A},\mathrm{T}}=0$ removes the constraint on the interpretation of triggered inputs leading to misaligned explanations in comparison to the original ones.
@@
             $\mathbf{x}^{\prime}\leftarrow\mathbf{x}\mathbin{\mathchoice{\ooalign{$\displaystyle\circ$\cr$\displaystyle\cdot$\cr}}{\ooalign{$\textstyle\circ$\cr$\textstyle\cdot$\cr}}{\ooalign{$\scriptstyle\circ$\cr$\scriptstyle\cdot$\cr}}{\ooalign{$\scriptscriptstyle\circ$\cr$\scriptscriptstyle\cdot$\cr}}}\mathbf{t}$
@@
             $\mathcal{L}_{\mathrm{adv}}\mathrel{+}=\lambda_{\mathrm{A},\neg\mathrm{T}}\mathcal{L}_{\mathrm{A},\neg\mathrm{T}}(\mathbf{z},\hat{\mathbf{z}})$
@@
             $\mathcal{L}_{\mathrm{adv}}\mathrel{+}=\lambda_{\mathrm{U},\neg\mathrm{T}}\mathcal{L}_{\mathrm{U},\neg\mathrm{T}}(\hat{\mathbf{p}})$
@@ 3.3 On the security of concept bottleneck models
The model can be trained end-to-end fashion by jointly optimizing loss on both class prediction and concept predictions. The original formulation considers an extension of the classification problem with $d_{\mathbf{p}}$ concepts where $\mathcal{D}_{\mathrm{train}}$ includes their labels, i.e. $\left\{\left(\mathbf{x}^{(1)},\mathbf{p}^{(1)},y^{(1)}\right),\ldots,\left(\mathbf{x}^{(n)},\mathbf{p}^{(n)},y^{(n)}\right)\right\}\in\mathcal{X}\times\mathcal{P}\times\mathcal{Y}$ and $\mathcal{P}\subseteq\{0,1\}^{d_{\mathbf{p}}}$.
As compared to ProtoViT described in Section 3.2, CBMs cannot be fooled with adversarial prototype (concept) manipulation due to a more strict training regime. Furthermore, backdoor attacks could be much easier to discover with the semantic labeled concepts that can be compared with the raw input representation, as compared to heatmaps in PIP-Net. Thus, more sophisticated approaches are needed for the manipulation of CBMs, e.g. by fooling concept embedding generators (Zarlenga et al., 2022), or concept generation and labeling (Yang et al., 2023; Oikarinen et al., 2023). We further discuss the differences between prototype-based networks and concept bottleneck models with regard to our adversarial analysis in Section 5.
@@ 4.1 Manipulating ProtoViT for bird species recognition
Following the original work of Ma et al. (2024), analyze ProtoViT based on three vision transformer backbones pre-trained on ImageNet (Russakovsky et al., 2015): CaiT-XXS-24 (Touvron et al., 2021), DeiT-Tiny & DeiT-Small (Touvron et al., 2021). We train these models on the CUB-200-201 (Birds, Wah et al., 2011) dataset using supplementary code and default hyperparameters. For prototype manipulation, we use two additional datasets: Stanford Cars (Cars, Krause et al., 2013) and a custom-made set of bird images of species not represented in Birds (Out-of-distribution Birds). Specifically, we collected 2044 images from the ImageNet training set from 13 bird species (classes), which we did not find in Birds. We first train each model for 30 epochs and then perform adversarial fine-tuning for 10 epochs. Each experiment is repeated 5 times to report the mean accuracy with standard deviations. Further details on datasets, preprocessing, and training hyperparameters are provided in Appendix A.
@@ Results.
Table 1 shows the test accuracy of all the trained and attacked models. On average, training with out-of-distribution birds leads to a drop in performance of 3.5% points accuracy, while for Cars, the performance drops significantly by about 12.5% points. This might be a reason why the prototype manipulativeness we discovered here was previously concealed. On average, our prototype manipulation improves the baseline of training with the new prototype set by 2% points on Out-of-distribution Birds and 9% points on Cars. The smallest and worst-performing model, DeiT-Tiny, is the most challenging to manipulate without a drop in performance. We provide more context on the stealthiness of prototype manipulation in Figure 4, which shows the decreasing test accuracy performance when substituting more prototypes, and the influence of fine-tuning the last classification layer after the substitution. These results emphasize the need to disentangle the influence of semantic similarity and classification layer weights in making interpretable predictions.
@@
Figure 4: Test accuracy performance as a function of the fraction of substituted prototypes for three vision transformer backbones: before and after finetuning the ProtoViT’s last layer.
@@
We further reinforce the claim that the predictive performance metric is not enough to evaluate the model’s interpretability. ProtoViT is able to achieve state-of-the-art performance using prototypes stemming from car images, e.g. CaiT-XXS-24 achieves 85.7% accuracy compared to 85.8% reported in the original work (Ma et al., 2024). Since we know that bird species and car brands are recognized by dissimilar semantic features, it is surprising that the network performs well using semantically dissimilar prototypes. It questions the inherent interpretability of the architecture understood as domain-specificity of prototypes.
@@
Is the prototype interpretable? Figure 5 displays a so-called global analysis of the DeiT-Small ProtoViT, where the nearest (sub)patches from the test set are shown for a representative set of prototypes (5 out of 2000) as their inherent interpretation. For context, we additionally display the nearest (sub)patches from Cars (on the right). Interestingly, in this case, prototype no. 488 maps black bird wings into black car wing doors, prototype no. 1020 maps bird belly into car headlight, no. 1619 maps orange color between birds and cars, and prototype no. 1995 seems to map a black bird’s head into the dim car window. Note that this intuition is only a biased approximation of what the model has actually learned. Such a visual confirmation process always occurs when interpreting prototype networks, even with perfect in-distribution images, i.e. the network architecture only looks like interpretable. User feedback claiming model correctness based on similarities between patches is not enough to evaluate model interpretability in this context because it is subject to human confirmation bias.
@@
Is the prediction interpretable? In Figure 6, we conduct a so-called local analysis of the same model, where the nearest or most contributing prototypes are shown for recognizing a few representative bird images. It seems like the model relies on color and spurious correlations for many of the predictions (Fig. 6B), e.g. it matches blue and red bird parts with blue and red car parts, respectively. The manipulated model only exaggerates the well-known issue of prototype redundancy, where the last classification layer contains interchangeable weights between correlated features (Fig. 6C). Lack of sparsity hinders interpretability. In general, complex, uninterpretable prototypes suppress our understanding of the model’s reasoning (Fig. 6D). It is further unclear whether one should pay more attention to the similarity score or classification weight when interpreting the prediction (Fig. 6E). We defer an analogous analysis of a model that uses out-of-distribution birds as prototypes to Appendix, Figure 10.
We perceive this high attack success rate demonstrates an inherent flaw in prototype-based networks, and it would generalize across other dataset pairs, specifically non-car prototypes, which is left for future work.
@@ 4.2 Backdoor attacks on PIP-Net for medical imaging
Instead of relying on benchmarks typical to evaluating prototype-based networks like Birds and Cars, we choose to evaluate backdoor attacks on a high-stakes application like medical image classification where security is of particular interest. Our motivation stems from an increasing interest in the application of these models in healthcare (Nauta et al., 2023a; Santi et al., 2024a; b).
@@
at the same time acknowledging that tuning these hyperparameters would have resulted in better results, which is out of the scope of this work. Each experiment is repeated 3 times to report the mean metric values with standard deviations. Further details on datasets, preprocessing, and training hyperparameters are provided in Appendix A.
@@
Dataset PIP-Net Backbone Accuracy ASR Alignment
@@
Dataset PIP-Net Backbone Accuracy ASR Alignment
@@
Metrics.
We measure the accuracy on the test set between the original and backdoored model, as well as the (mis)alignment loss for the original and triggered inputs between both models. We report the attack success rate as the fraction of triggered inputs leading to changed predictions.
Quantitative results.
Tables 3 & 3 report predictive performance, attack success rate, and (mis)alignment values for the disguising and red herring attacks, respectively. The attack fine-tuning can result in up to 5% points accuracy drop in our setup. We maintain about 85% attack success rate, which is high considering a multi-trigger multi-place threat model. In general, it is considerably harder to successfully embed the backdoor in models trained on MURA, especially in ResNets, which we naturally attribute to the fact that it is harder to train models on MURA (than ISIC) and ResNets (than newer ConvNeXt) in the first place.
Note that were not able to reproduce the out-of-distribution detection results as reported in (Nauta et al., 2023b, table 2). In our setup, the models predicted that (on average, about) 65% inputs from Birds are in distribution for both ISIC and MURA, even under more favorable thresholds than the proposed 95%. Crucially, the models did not ever “abstain” from the decision made for out-of-distribution inputs as intended. It further emphasizes the need for more robust defense mechanisms built into (interpretable) deep learning architectures.
Regarding the optimized (mis)alignment of explanations for inputs with and without the trigger, it might be hard to compare them between datasets and models due to not-obvious normalization. On average, there is only a relative 1.3×1.3$\times$1.3 × increase in misalignment for the disguising scenario, compared to 26×26$\times$26 × for the red herring, which properly lines up with the defined training objective. In all cases, alignment values for inputs with a trigger in the disguising scenario are higher than those without a trigger in the red herring scenario. There definitely is a trade-off between maintaining high backdoored model accuracy and alignment with the original model, as well as the expected attack success rate for a given use case. All in all, we find the ConvNeXt architecture considerably easier to attack, as evidenced by all the metric values.
@@
A subtle yet obvious realization after attacking PIP-Net is that a global analysis of a prototype-based network cannot immediately show whether a backdoor exists because the prototype projection set (or visualization set in the case of PIP-Net) contains no images with triggers (we defer such example to Appendix, Figure 12). Hence, alternative defense mechanisms are needed.

### v2

@@ 1 $\mathbf{Z},\mathbf{P},\mathbf{I},\hat{\theta}_{g}\leftarrow[\;],[\;],[\;],[\;]$
In what follows, let $\mathcal{D}_{\mathrm{OOD}}$ denote a set of inputs an adversary chooses for prototype manipulation. The naive approach is to only substitute $\mathcal{D}_{\mathrm{train}}$ with $\mathcal{D}_{\mathrm{OOD}}$ for the single prototype projection stage during training. Specifically for ProtoViT, we propose to randomly assign inputs in $\mathcal{D}_{\mathrm{OOD}}$ to the original labels from $\mathcal{Y}$ to satisfy the additional constraint mentioned above. We found this baseline to achieve subpar results (as later shown in Table 1) and thus implemented a simple matching algorithm instead. We substitute latent representations stored in $\theta_{g}$ with the closest embeddings of parts of inputs in $\mathcal{D}_{\mathrm{OOD}}$ as denoted with $\hat{\theta}_{g}$ and given in Algorithm 1. In practice, iterating over the prototype projection set (Alg. 1, line 2) is executed in batches similar to training. Note that prototype substitution can be done at any point, e.g. after the first stage of training encoder parameters $\theta_{f}$, or after the last stage of training classification parameters $\theta_{h}$ when parameters $\theta_{f}$ are frozen. In general, substituting prototypes introduces an approximation error $\theta_{g}-\hat{\theta}_{g}||$ that propagates to errors in predictions of $g$ and finally $h$. One can correct the predictions of $h$ by fine-tuning parameters $\theta_{h}$, which we further analyze in Section 5.1.
@@
4.2 Backdoor attacks
The goal of a backdoor attack is to change the model’s prediction for inputs with a specific trigger. It exploits the well-known problem that deep neural networks solve problems by taking shortcuts instead of learning the intended solution (Geirhos et al., 2020). Such an attack could be discovered with post-hoc explanation methods (Noppel et al., 2023). Thus, a successful disguising backdoor attack in the context of interpretability assumes concealing the trigger in a visualization of the model’s explanation. Figure 2 shows an intuition behind our attack on interpretable deep learning models. In an extreme scenario, an adversary might aim to exaggerate the evidence for a manipulated prediction so as to distract from the attack, aka red herring.
Note that natural semantic triggers already exist in data in the form of bias, e.g. commonly in medical images (Nauta et al., 2023a). Figure 3 shows features occurring coincidentally with the predicted outcome while having no causal influence on it (see Mikołajczyk et al., 2022, and references given there for a discussion on this issue). For example, hair appearing in skin images or markers appearing in X-rays are shortcuts used by neural networks to achieve better in-training performance, diminishing their generalization. Although we implement the trigger as a small adversarial patch in an image corner to observe the qualitative result, our method is general in nature. Intuitively, such a backdoor could work because there exists no prototype to be matched with a trigger, so the model chooses another prototype to match instead. However, as we show in Section 5.1, the model is able to learn superficial prototypes with Algorithm 1 if needed.
@@ 4.2 Backdoor attacks
[…]
$\mathcal{L}_{\mathrm{adv}}(\hat{\theta})=\lambda_{\mathrm{C}}\mathcal{L}_{\mathrm{C}}+\lambda_{\mathrm{A},\neg\mathrm{T}}\mathcal{L}_{\mathrm{A},\neg\mathrm{T}}+\lambda_{\mathrm{A},\mathrm{T}}\mathcal{L}_{\mathrm{A},\mathrm{T}}+\lambda_{\mathrm{U},\neg\mathrm{T}}\mathcal{L}_{\mathrm{U},\neg\mathrm{T}}+\lambda_{\mathrm{U},\mathrm{T}}\mathcal{L}_{\mathrm{U},\mathrm{T}},$ (5)
where $\mathcal{L}_{\mathrm{A}}$ & $\mathcal{L}_{\mathrm{U}}$ follow Equations 2 & 3, respectively, $\mathrm{T}$ denotes operations on embeddings of inputs from $\mathcal{D}_{\mathrm{trigger}}$ and $\neg\mathrm{T}$ on the original ones. Algorithm 2 describes in detail the exemplary computation of a backdoor attack for binary classification with trigger $\mathbf{t}$, where iterating over the training set (Alg. 2, line 3) is executed in batches in practice. Our implementation greatly benefits from using 2 GPUs to store model parameters, inputs, and outputs between $\theta$ and $\hat{\theta}$. Although the poisoning ratio of the dataset is an arbitrary hyperparameter of the method, we found in experiments that poisoning the whole dataset and fine-tuning for over a single epoch is sufficient to successfully embed the backdoor. Manipulating $\lambda_{\mathrm{A},\neg\mathrm{T}}$ and $\lambda_{\mathrm{A},\mathrm{T}}$ hyperparameters produces a spectrum of results between disguising and red herring backdoor attacks. For example, setting $\lambda_{\mathrm{A},\mathrm{T}}=0$ removes the constraint on the interpretation of triggered inputs, leading to misaligned explanations in comparison to the original ones.
@@
2for $\mathrm{epoch}\;\mathbf{in}\;1\ldots\mathrm{epochs}$ do
@@
[…]
@@
[…]
@@ 4.3 On the security of concept bottleneck models
The model can be trained in an end-to-end fashion by jointly optimizing the loss on both class prediction and concept predictions. The original formulation considers an extension of the classification problem with $d_{\mathbf{p}}$ concepts where $\mathcal{D}_{\mathrm{train}}$ includes their labels, i.e. $\left\{\left(\mathbf{x}^{(1)},\mathbf{p}^{(1)},y^{(1)}\right),\ldots,\left(\mathbf{x}^{(n)},\mathbf{p}^{(n)},y^{(n)}\right)\right\}\in\mathcal{X}\times\mathcal{P}\times\mathcal{Y}$ and $\mathcal{P}\subseteq\{0,1\}^{d_{\mathbf{p}}}$.
As compared to ProtoViT described in Section 2, CBMs cannot be fooled with adversarial prototype (concept) manipulation due to a stricter training regime. Furthermore, backdoor attacks could be much easier to discover with the semantically labeled concepts that can be compared with the raw input representation, as compared to heatmaps in PIP-Net. Thus, more sophisticated approaches are needed for the manipulation of CBMs, e.g. by fooling concept embedding generators (Zarlenga et al., 2022), or concept generation and labeling (Yang et al., 2023; Oikarinen et al., 2023). We further discuss the differences between prototype-based networks and concept bottleneck models with regard to our adversarial analysis in Section 6.
@@ 5.1 Manipulating ProtoViT for bird species recognition
Following the original work of Ma et al. (2024), analyze ProtoViT based on three vision transformer backbones pre-trained on ImageNet (Russakovsky et al., 2015): CaiT-XXS-24 (Touvron et al., 2021), DeiT-Tiny & DeiT-Small (Touvron et al., 2021). We train these models on the CUB-200-2011 (Birds, Wah et al., 2011) dataset using supplementary code and default hyperparameters. For prototype manipulation, we use two additional datasets: Stanford Cars (Cars, Krause et al., 2013) and a custom-made set of bird images of species not represented in Birds (Out-of-distribution Birds). Specifically, we collected 2044 images from the ImageNet training set from 13 bird species (classes), which we did not find in Birds. We first train each model for 30 epochs and then perform adversarial fine-tuning for 10 epochs. Each experiment is repeated 5 times to report the mean accuracy with standard deviations. Further details on datasets, preprocessing, and training hyperparameters are provided in Appendix A.
@@ Results.
Table 1 shows the test accuracy of all the trained and attacked models. On average, training with out-of-distribution birds leads to a drop in accuracy performance of 3.5% points, while for Cars, the performance drops significantly by about 12.5% points. This might be a reason why the prototype manipulativeness we discovered here was previously concealed. On average, our prototype manipulation improves the baseline of training with the new prototype set by 2% points on Out-of-distribution Birds and 9% points on Cars. The smallest and worst-performing model, DeiT-Tiny, is the most challenging to manipulate without a drop in performance. We provide more context on the stealthiness of prototype manipulation in Figure 4, which shows the decreasing test accuracy performance when substituting more prototypes, and the influence of fine-tuning the last classification layer after the substitution. These results emphasize the need to disentangle the influence of semantic similarity and classification layer weights in making interpretable predictions.
@@
Figure 4: Test accuracy performance as a function of the fraction of substituted prototypes for three vision transformer backbones: before and after fine-tuning the ProtoViT’s last layer.
@@
We further reinforce the claim that the predictive performance metric is not enough to evaluate the model’s interpretability. ProtoViT is able to achieve state-of-the-art performance using prototypes stemming from car images, e.g. CaiT-XXS-24 achieves 85.7% accuracy compared to 85.8% reported in the original work (Ma et al., 2024). Since we know that bird species and car brands are recognized by dissimilar semantic features, it is surprising that the network performs well using semantically dissimilar prototypes. It questions the inherent interpretability of the architecture understood as the domain-specificity of prototypes.
@@
Is the prototype interpretable?
Figure 5 displays a so-called global analysis of the DeiT-Small ProtoViT, where the nearest (sub)patches from the test set are shown for a representative set of prototypes (5 out of 2000) as their inherent interpretation. For context, we additionally display the nearest (sub)patches from Cars (on the right). Interestingly, in this case, prototype no. 488 maps black bird wings into black car wing doors, prototype no. 1020 maps the bird’s belly into the car’s headlight, no. 1619 maps the orange color between birds and cars, and the prototype no. 1995 seems to map a black bird’s head into the dim car window. Note that this intuition is only a biased approximation of what the model has actually learned. Such a visual confirmation process always occurs when interpreting prototype networks, even with perfect in-distribution images, i.e. the network architecture only looks like interpretable. User feedback claiming model correctness based on similarities between patches is not enough to evaluate model interpretability in this context because it is subject to human confirmation bias.
@@
Is the prediction interpretable? In Figure 6, we conduct a so-called local analysis of the same model, where the nearest or most contributing prototypes are shown for recognizing a few representative bird images. It seems like the model relies on color and spurious correlations for many of the predictions (Fig. 6B), e.g. it matches blue and red bird parts with blue and red car parts, respectively. The manipulated model only exaggerates the well-known issue of prototype redundancy, where the last classification layer contains interchangeable weights between correlated features (Fig. 6C). Lack of sparsity hinders interpretability. In general, complex, uninterpretable prototypes suppress our understanding of the model’s reasoning (Fig. 6D). It is further unclear whether one should pay more attention to the similarity score or classification weight when interpreting the prediction (Fig. 6E). We defer an analogous analysis of a model that uses out-of-distribution birds as prototypes to Appendix, Figure 10. We perceive that such a high attack success rate demonstrates an inherent flaw in prototype-based networks, and it would generalize across other dataset pairs, specifically non-car prototypes, which is left for future work.
@@ 5.2 Backdoor attacks on PIP-Net for medical imaging
Instead of relying on benchmarks typical of evaluating prototype-based networks like Birds and Cars, we choose to evaluate backdoor attacks on a high-stakes application like medical image classification, where security is of particular interest. Our motivation stems from an increasing interest in the application of these models in healthcare (Nauta et al., 2023a; Santi et al., 2024a; b).
@@
at the same time, acknowledging that tuning these hyperparameters would have resulted in better results, which is out of the scope of this work. We backdoor the whole dataset as shown in Algorithm 2 and fine-tune the model for 3 epochs. Each experiment is repeated 3 times to report the mean metric values with standard deviations. Further details on datasets, preprocessing, and training hyperparameters are provided in Appendix A.
Metrics.
We measure the accuracy on the test set between the original and backdoored model, as well as the (mis)alignment loss for the original ($\mathcal{L}_{\mathrm{A},\neg\mathrm{T}}$) and triggered ($\mathcal{L}_{\mathrm{A},\mathrm{T}}$) inputs between both models. Refer to Appendix A.1 for details on the class balance. We report the attack success rate as the fraction of triggered inputs leading to changed predictions.
Quantitative results.
Tables 3 & 3 report predictive performance, attack success rate, and (mis)alignment values for the disguising and red herring attacks, respectively. The attack fine-tuning can result in up to 5% points of accuracy drop in our setup. We maintain about an 85% attack success rate, which is high considering a multi-trigger multi-place threat model. In general, it is considerably harder to successfully embed the backdoor in models trained on MURA, especially in ResNets, which we naturally attribute to the fact that it is harder to train models on MURA (than ISIC) and ResNets (than newer ConvNeXt) in the first place.
@@
Dataset PIP-Net Backbone Accuracy ASR Alignment ($\mathcal{L}_{\mathrm{A},\cdot}$)
@@
Dataset PIP-Net Backbone Accuracy ASR Alignment ($\mathcal{L}_{\mathrm{A},\cdot}$)
@@
To give a broader context, we apply the PIP-Net’s out-of-distribution detection method as reported in (Nauta et al., 2023b, table 2). In our setup, the models predicted that (on average, about) 65% of inputs from Birds are in distribution for both ISIC and MURA, even under more favorable thresholds than the proposed 95%. Crucially, the models never “abstained” from the decision made for out-of-distribution inputs as intended. It further emphasizes the need for more robust defense mechanisms built into (interpretable) deep learning architectures.
Regarding the optimized (mis)alignment of explanations for inputs with and without the trigger, it might be hard to compare them between datasets and models due to non-obvious normalization. On average, there is only a relative $1.3\times$ increase in misalignment for the disguising scenario, compared to $26\times$ for the red herring, which properly lines up with the defined training objective. In all cases, alignment values for inputs with a trigger in the disguising scenario are higher than those without a trigger in the red herring scenario. There definitely is a trade-off between maintaining high backdoored model accuracy and alignment with the original model, as well as the expected attack success rate for a given use case. All in all, we find the ConvNeXt architecture considerably easier to attack, as evidenced by all the metric values.
@@
A subtle yet obvious realization after attacking PIP-Net is that a global analysis of a prototype-based network cannot immediately show whether a backdoor exists because the prototype projection set (or visualization set in the case of PIP-Net) contains no images with triggers (we defer such an example to Appendix, Figure 12). Hence, alternative defense mechanisms are needed.